(b) Suppose you wish to invest $2,000 today so that you have $4,000 six years from
now. What must the annual interest rate be in order to achieve your goal if interest is
compounded quarterly? (3 marks)
Answer #1

Ans 11.72%

The annual interest rate be in order to achieve your goal if interest is compounded quarterly is 11.72%

FV = Future Value PV = Present Value r= rate of interest n= no of period FV/ PV = (1+r%)^n 4000 / 2000 = (1+r%/4)^24 2 = (1 +